1. find the domains of f(x) =


√x+7 3

______  and g(x) =   _______ 

 x+4   |x| + 2



 

            

                 

          

 


1
Expert's answer
2021-10-06T18:02:12-0400
"f(x)=\dfrac{\sqrt{x+7}}{3}"

"x+7\geq0=>x\geq-7"

"Domain: [-7,\infin)"



"g(x)=\dfrac{x+4}{|x|+2}"

"|x|\not=-2=>x\in \R"

"Domain: (-\infin,\infin)"